### v2

@@ 1 Introduction
Suppose the graph $G$ has $n$ vertices and $m$ edges. Computing an individual connectivity value in $G$ is easy: since the maximum flow between a fixed pair of nodes can be found in almost linear time [CKL+22], we know that for any fixed pair of vertices $(s,t)$ in $G$, we can compute $\lambda(s,t)$ in essentially optimal $m^{1+o(1)}$ time. For many applications however, knowing a single connectivity is not so useful, and it is instead far more informative to know the values of multiple connectivities in a graph.
This motivates the All-Pairs Connectivity (APC) problem, where we are tasked with computing connectivities for all pairs of vertices in a given graph. A long line of work recently culminated in a near-optimal $\tilde{O}(n^{2})$ time algorithm for solving APC over undirected graphs [AKL+21]. Throughout this paper, we focus on the general case where $G$ is directed.
@@
The $k$-APC problem is relevant in contexts where knowing the precise connectivity values between “well-connected” nodes is not important, and instead we care more about distinguishing for each pair of vertices whether its connectivity is small or large (where $k$ is our cutoff for what counts as “small” and “large”). Since the connectivity between any pair of nodes in $G$ is at most $n-1$, $k$-APC recovers the general APC problem when $k=n-1$. As $k$ gets smaller, $k$-APC intuitively becomes easier. Indeed, it was recently shown that for any positive integer $k$, $k$-APC can be solved in $\tilde{O}((kn)^{\omega})$ time [AJ23].
@@ 2 Preliminaries
For any edge $e=(u,v)\in E$, we let $\operatorname{tail}(e)=u$ denote the node $e$ exits, and $\operatorname{head}(e)=v$ denote the node $e$ enters.
@@
Given a matrix $M$, for any row index $i$ and column index $j$ we let $M[i,j]$ be the $(i,j)$ entry of $M$. Given subsets $I$ and $J$ of row and column indices respectively, we let $M[I,J]$ be the submatrix of $M$ restricted to rows in $I$ and columns in $J$. We also let $M[I,\cdot]$ be the submatrix restricted to rows in $I$ and all columns, and $M[\cdot,J]$ be the submatrix on all rows and restricted to columns in $J$. We let $\operatorname{rank}M$ denote the rank of $M$, defined to be largest nonnegative integer $r$ such that $M$ contains an $r\times r$ submatrix with nonzero determinant. When $M$ is a square matrix, we let $\det M$ denote the determinant of $M$, $\operatorname{adj}(M)$ denote the adjoint of $M$, and $M^{-1}$ denote the inverse of $M$ (if it exists).
@@ 3 Power Series Preliminaries
If $F$ is a formal power series with constant term $1$, then there is a unique formal series $H=F^{-1}$ with constant term $1$ satisfying $F\cdot H=1$.
@@
Suppose $X$ is a square matrix with polynomial entries such that every entry of $X$ has zero constant term. Then
@@
with entries restricted to terms of degree at most $d$. Since every nonzero entry of $X^{d+1}$ has degree greater than $d$, we see that $M_{d}=I$ is the identity matrix. Since this equation holds for every fixed $d\geq 0$, Equation 4 holds as well.
@@ 4 Enumeration with Edge-Adjacency Matrices
We enumerate walks not by counting their number, but by assigning each walk a certain monomial weight, which records information about the edges traversed in the walk. Enumeration for our purposes corresponds to summing the weights of all walks (or collections of walks) in a certain family of interest.
@@
$\omega(W)=\prod_{j=1}^{\ell-1}x_{e_{j}e_{j+1}}$
of $W$ be the monomial $\omega(W)$ recording all pairs of consecutive edges traversed by $W$. By convention, we assign a walk $W$ of length one (i.e., a single edge) weight $\omega(W)=1$. More generally, given a collection of walks $\mathcal{C}=\langle W_{1},\dots,W_{r}\rangle$ we let the weight
$\omega(\mathcal{C})=\prod_{j=1}^{r}\omega(W_{j})$
@@
$(I-X)^{-1}[e,f]=\sum_{\ell=0}^{\infty}\left(\sum_{W\in\mathcal{W}_{\ell+1}(e,f)}\omega(W)\right).$
@@
$\det\leavevmode\nobreak\ (I-X)^{-1}[S,T]=\sum_{\ell=1}^{\infty}\left(\sum_{\mathcal{C}\in\mathcal{F}_{\ell}(S,T)}\omega(\mathcal{C})\right)$
@@
$\Gamma[e,\pi(e)]=\sum_{\ell=0}^{\infty}\left(\sum_{W\in\mathcal{W}_{\ell+1}(e,\pi(e))}\omega(W)\right).$ (7)
@@
$\prod_{e\in S}\Gamma[e,\pi(e)]=\prod_{i=1}^{r}\left(\sum_{\ell=0}^{\infty}\left(\sum_{W\in\mathcal{W}_{\ell+1}(e_{i},\pi(e_{i}))}\omega(W)\right)\right).$ (8)
@@
$\prod_{i=1}^{r}\left(\sum_{\ell=0}^{\infty}\left(\sum_{W\in\mathcal{W}_{\ell+1}(e_{i},\pi(e_{i}))}\omega(W)\right)\right)=\sum_{\ell=1}^{\infty}\left(\sum_{\begin{subarray}{c}(\ell_{1},\dots,\ell_{r})\in\mathcal{L}\\ W_{i}\in\mathcal{W}_{\ell_{i}}(e_{i},\pi(e_{i}))\end{subarray}}\prod_{i=1}^{r}\omega(W_{i})\right).$
@@
$\det\leavevmode\nobreak\ \Gamma[S,T]=\sum_{\ell=1}^{\infty}\left(\sum_{\pi\in\mathfrak{S}(S,T)}\sum_{\begin{subarray}{c}\\ (\ell_{1},\dots,\ell_{r})\in\mathcal{L}\\ W_{i}\in\mathcal{W}_{\ell_{i}}(e_{i},\pi(e_{i}))\end{subarray}}\prod_{i=1}^{r}\omega(W_{i})\right).$ (9)
@@
$\omega(\mathcal{C})=\prod_{i=1}^{r}\omega(W_{i}),$
@@
$\det\Gamma[S,T]=\sum_{\ell=1}^{\infty}\left(\sum_{\mathcal{C}\in\mathcal{F}_{\ell}(S,T)}\omega(\mathcal{C})\right)$
@@
$\sum_{\mathcal{C}\in\mathcal{F}_{\ell}(S,T)}\omega(\mathcal{C})=\sum_{\mathcal{C}\in\mathcal{D}_{\ell}(S,T)}\omega(\mathcal{C}).$
@@
$\sum_{\mathcal{C}\in\mathcal{F}}\omega(\mathcal{C})=\sum_{\mathcal{C}\in\mathcal{D}}\omega(\mathcal{C})+\sum_{\mathcal{C}\in\mathcal{S}}\omega(\mathcal{C}).$
@@
$\sum_{\mathcal{C}\in\mathcal{S}}\omega(\mathcal{C})$
is the zero polynomial. We prove this by pairing up collections $\mathcal{C}$ in $\mathcal{S}$ of equal weight $\omega(\mathcal{C})$, and observing that contributions from such collections vanish modulo two.
@@
Figure 1: Given walks $W_{i}=\langle e_{1},e_{2},g,e_{3},e_{4}\rangle$ and $W_{j}=\langle f_{1},f_{2},g,f_{3},f_{4}\rangle$ overlapping at $g$, we can swap their suffixes to produce walks $W^{\prime}_{i}=\langle e_{1},e_{2},g,f_{3},f_{4}\rangle$ and $W^{\prime}_{j}=\langle f_{1},f_{2},g,e_{3},e_{4}\rangle$ which still overlap at $g$. The weights for both pairs $\omega(W_{i},W_{j})=(x_{e_{1}e_{2}}x_{e_{2}g}x_{ge_{3}}x_{e_{3}e_{4}})\cdot(x_{f_{1}f_{2}}x_{f_{2}g}x_{gf_{3}}x_{f_{3}f_{4}})$ and $\omega(W^{\prime}_{i},W^{\prime}_{j})=(x_{e_{1}e_{2}}x_{e_{2}g}x_{gf_{3}}x_{f_{3}f_{4}})\cdot(x_{f_{1}f_{2}}x_{f_{2}g}x_{ge_{3}}x_{e_{3}e_{4}})$ agree because each pair traverses the same multiset of consecutive pairs of edges.
@@
as the concatenation of a prefix walk $A_{i}$ not including edge $e$, and a suffix walk $B_{i}$ which begins with edge $e$. We can similarly split $W_{j}$ uniquely
@@
Indeed, for all $l\in[r]$ with $l<i$, the walk $W^{\prime}_{l}=W_{l}$ does not intersect any other walk in $\mathcal{C}$ at an edge, by the definition of $i$. Since the multiset of edges traversed by walks in $\mathcal{C}\setminus\left\{W_{l}\right\}$ and $\mathcal{C}\setminus\left\{W^{\prime}_{l}\right\}$ are the same, this means that $W^{\prime}_{l}$ does not intersect any other walk in $\mathcal{C}^{\prime}$ at an edge either. So $i$ is the smallest index in $[r]$ such that $W^{\prime}_{i}$ intersects some other walk in $\mathcal{C}^{\prime}$ at an edge. Since the prefixes of $W^{\prime}_{i}$ and $W_{i}$ before edge $e$ are the same, we see that $e$ is also the first edge in $W^{\prime}_{i}$ which is contained in another walk of $\mathcal{C}^{\prime}$. Then because $W_{j}^{\prime}$ traverses edge $e$, and $W_{l}^{\prime}=W_{l}$ for all $l\not\in\left\{i,j\right\}$, we get that $j>i$ is the smallest index such that $W_{j}^{\prime}$ overlaps with $W^{\prime}_{i}$ at edge $e$. Then when we swap the suffixes of $W^{\prime}_{i}$ and $W^{\prime}_{j}$ after the first appearance of $e$ on these walks, we recover $W_{i}$ and $W_{j}$ respectively, and so applying the suffix swapping procedure to $\mathcal{C}^{\prime}$ produces the original collection $\mathcal{C}$ as claimed.
@@
$\omega(\mathcal{C})=\omega(\mathcal{C}^{\prime})$
@@
$\omega(\mathcal{C})+\omega(\mathcal{C}^{\prime})=0.$
@@
$\sum_{\mathcal{C}\in\mathcal{S}}\omega(\mathcal{C})=0.$
@@
In the proof of Lemma 4.7, our argument used the fact that we work over a field of characteristic two. This restriction on the characteristic is only included for the sake of simplicity, and is not necessary for the result to hold. Indeed, if we instead worked over a field of odd characteristic, then all that changes is that terms in the expansion of the determinant come with a sign, and we can now pair up and cancel terms with opposite signs.
@@
$\det\leavevmode\nobreak\ (I-X)^{-1}[S,T]=\sum_{\ell=1}^{\infty}\left(\sum_{\mathcal{C}\in\mathcal{D}_{\ell}(S,T)}\omega(\mathcal{C})\right).$
@@
Suppose $\mathcal{C}=\langle P_{1},\dots,P_{r}\rangle$ is a collection of $r$ edge-disjoint paths from $S$ to $T$ in $G$. Then the term $\omega(\mathcal{C})$ occurs in the expansion of
@@
given by Corollary 4.10. Moreover, any collection of paths $\mathcal{C}^{\prime}\neq\mathcal{C}$ from $S$ to $T$ has weight $\omega(\mathcal{C}^{\prime})\neq\omega(\mathcal{C})$, because $\mathcal{C}$ consists of edge-disjoint paths (so looking at the variables appearing in $\omega(\mathcal{C})$, we can recover $\mathcal{C}$ uniquely). Hence, no other term from the summation in Corollary 4.10 produces the same monomial $\omega(\mathcal{C})$. So $\omega(\mathcal{C})$ appears in Equation 10 with nonzero coefficient, which implies that the determinant from Equation 10 is a nonzero formal power series.
@@
DAGs are an interesting case for APC and $k$-APC, since we do not know of algorithms solving these problems on DAGs faster than the case of general directed graphs, and the best conditional lower bounds we have for these problems hold in the case of DAGs [AGI+18].
@@
First, any walk in a DAG has length at most $n-1$. Hence over DAGs, $X^{\ell}$ is the all-zeros matrix for $\ell\geq n$, so Proposition 3.3 shows that $\Gamma=(I-X)^{-1}$ is a polynomial matrix. Thus, determinants of submatrices of $\Gamma$ are polynomials instead of formal power series.
Second, any walk in a DAG is a path. Thus we can skip the step in Lemma 4.12 of going from edge-disjoint walks to edge-disjoint paths.

### v3

@@ 1 Introduction
Suppose the graph $G$ has $n$ vertices and $m$ edges. Computing an individual connectivity value in $G$ is easy: since the maximum flow between a fixed pair of nodes can be found in almost linear time [CKL+23], we know that for any fixed pair of vertices $(s,t)$ in $G$, we can compute $\lambda(s,t)$ in almost-optimal $m^{1+o(1)}$ time. For many applications however, knowing a single connectivity is not so useful, and it is instead far more informative to know the values of multiple connectivities in a graph.
This motivates the All-Pairs Connectivity (APC) problem, where we are tasked with computing connectivities for all pairs of vertices in a given graph. A long line of work recently culminated in a near-optimal $\tilde{O}(n^{2})$ time algorithm for solving APC over undirected graphs [AKL+22]. Throughout this paper, we focus on the general case where $G$ is directed.
@@
The $k$-APC problem is relevant in contexts where knowing the precise connectivity values between “well-connected” nodes is not important, and instead we care more about distinguishing for each pair of vertices whether its connectivity is small or large (where $k$ is our cutoff for what counts as “small” and “large”). Since the connectivity between any pair of nodes in $G$ is at most $n-1$, the $k$-APC problem recovers the general APC problem when $k=n-1$. As $k$ gets smaller, $k$-APC intuitively becomes easier. Indeed, it was recently shown that for any positive integer $k$, $k$-APC can be solved in $\tilde{O}((kn)^{\omega})$ time [AJ23].
@@ 2 Preliminaries
For a directed edge $e=(u,v)\in E$ going from $u$ to $v$, we let $\operatorname{tail}(e)=u$ and $\operatorname{head}(e)=v$ denote the vertices that $e$ exits and enters respectively.
@@
Given a matrix $M$, for any row index $i$ and column index $j$ we let $M[i,j]$ be the $(i,j)$ entry of $M$. Given subsets $I$ and $J$ of row and column indices respectively, we let $M[I,J]$ be the submatrix of $M$ restricted to rows in $I$ and columns in $J$. We also let $M[I,\cdot]$ be the submatrix restricted to rows in $I$ and all columns, and $M[\cdot,J]$ be the submatrix on all rows and restricted to columns in $J$. We let $\operatorname{rank}M$ denote the rank of $M$, defined to be largest nonnegative integer $r$ such that $M$ contains an $r\times r$ submatrix with nonzero determinant. When $M$ is a square matrix, we let $\det M$ denote the determinant of $M$, $\operatorname{adj}(M)$ denote the adjugate of $M$, and $M^{-1}$ denote the inverse of $M$ (if it exists).
@@ 3 Power Series Preliminaries
If $F$ is a formal power series with constant term $1$, then there is a unique formal series $H=F^{-1}$ such that $F\cdot H=1$. Moreover, the constant term of $H$ equals $1$.
@@
Suppose $X$ is a square matrix with polynomial entries such that every entry of $X$ has constant term equal to zero. Then
@@
with entries restricted to terms of degree at most $d$. Since every nonzero entry of $X^{d+1}$ has degree greater than $d$, we see that $M_{d}=I$ is the identity matrix. Since this equation holds for every fixed $d\geq 0$, the product in Equation 4 equals the identity matrix as claimed.
@@ 4 Enumeration with Edge-Adjacency Matrices
We enumerate walks not by counting their number, but by assigning each walk a monomial weight, that records information about the edges traversed in the walk. Enumeration for our purposes corresponds to summing the weights of all walks (or collections of walks) in a certain family of interest.
@@
$\xi(W)=\prod_{j=1}^{\ell-1}x_{e_{j}e_{j+1}}$
of $W$ be the monomial $\xi(W)$ recording all pairs of consecutive edges traversed by $W$. By convention, we assign a walk $W$ of length one (i.e., a single edge) weight $\xi(W)=1$. More generally, given a collection of walks $\mathcal{C}=\langle W_{1},\dots,W_{r}\rangle$ we let the weight
$\xi(\mathcal{C})=\prod_{j=1}^{r}\xi(W_{j})$
@@
$(I-X)^{-1}[e,f]=\sum_{\ell=0}^{\infty}\left(\sum_{W\in\mathcal{W}_{\ell+1}(e,f)}\xi(W)\right).$
@@
$\det\leavevmode\nobreak\ (I-X)^{-1}[S,T]=\sum_{\ell=1}^{\infty}\left(\sum_{\mathcal{C}\in\mathcal{F}_{\ell}(S,T)}\xi(\mathcal{C})\right)$
@@
$\Gamma[e,\pi(e)]=\sum_{\ell=0}^{\infty}\left(\sum_{W\in\mathcal{W}_{\ell+1}(e,\pi(e))}\xi(W)\right).$ (7)
@@
$\prod_{e\in S}\Gamma[e,\pi(e)]=\prod_{i=1}^{r}\left(\sum_{\ell=0}^{\infty}\left(\sum_{W\in\mathcal{W}_{\ell+1}(e_{i},\pi(e_{i}))}\xi(W)\right)\right).$ (8)
@@
$\prod_{i=1}^{r}\left(\sum_{\ell=0}^{\infty}\left(\sum_{W\in\mathcal{W}_{\ell+1}(e_{i},\pi(e_{i}))}\xi(W)\right)\right)=\sum_{\ell=1}^{\infty}\left(\sum_{\begin{subarray}{c}(\ell_{1},\dots,\ell_{r})\in\mathcal{L}\\ W_{i}\in\mathcal{W}_{\ell_{i}}(e_{i},\pi(e_{i}))\end{subarray}}\prod_{i=1}^{r}\xi(W_{i})\right).$
@@
$\det\leavevmode\nobreak\ \Gamma[S,T]=\sum_{\ell=1}^{\infty}\left(\sum_{\pi\in\mathfrak{S}(S,T)}\sum_{\begin{subarray}{c}\\ (\ell_{1},\dots,\ell_{r})\in\mathcal{L}\\ W_{i}\in\mathcal{W}_{\ell_{i}}(e_{i},\pi(e_{i}))\end{subarray}}\prod_{i=1}^{r}\xi(W_{i})\right).$ (9)
@@
$\xi(\mathcal{C})=\prod_{i=1}^{r}\xi(W_{i}),$
@@
$\det\Gamma[S,T]=\sum_{\ell=1}^{\infty}\left(\sum_{\mathcal{C}\in\mathcal{F}_{\ell}(S,T)}\xi(\mathcal{C})\right)$
@@
$\sum_{\mathcal{C}\in\mathcal{F}_{\ell}(S,T)}\xi(\mathcal{C})=\sum_{\mathcal{C}\in\mathcal{D}_{\ell}(S,T)}\xi(\mathcal{C}).$
@@
$\sum_{\mathcal{C}\in\mathcal{F}}\xi(\mathcal{C})=\sum_{\mathcal{C}\in\mathcal{D}}\xi(\mathcal{C})+\sum_{\mathcal{C}\in\mathcal{S}}\xi(\mathcal{C}).$
@@
$\sum_{\mathcal{C}\in\mathcal{S}}\xi(\mathcal{C})$
is the zero polynomial. We prove this by pairing up collections $\mathcal{C}$ in $\mathcal{S}$ of equal weight $\xi(\mathcal{C})$, and observing that contributions from such collections vanish modulo two.
@@
Figure 1: Given walks $W_{i}=\langle e_{1},e_{2},g,e_{3},e_{4}\rangle$ and $W_{j}=\langle f_{1},f_{2},g,f_{3},f_{4}\rangle$ overlapping at $g$, we can swap their suffixes to produce walks $W^{\prime}_{i}=\langle e_{1},e_{2},g,f_{3},f_{4}\rangle$ and $W^{\prime}_{j}=\langle f_{1},f_{2},g,e_{3},e_{4}\rangle$ which still overlap at $g$. The weights for both pairs $\xi(W_{i},W_{j})=(x_{e_{1}e_{2}}x_{e_{2}g}x_{ge_{3}}x_{e_{3}e_{4}})\cdot(x_{f_{1}f_{2}}x_{f_{2}g}x_{gf_{3}}x_{f_{3}f_{4}})$ and $\xi(W^{\prime}_{i},W^{\prime}_{j})=(x_{e_{1}e_{2}}x_{e_{2}g}x_{gf_{3}}x_{f_{3}f_{4}})\cdot(x_{f_{1}f_{2}}x_{f_{2}g}x_{ge_{3}}x_{e_{3}e_{4}})$ agree because each pair traverses the same multiset of consecutive pairs of edges.
@@
as the concatenation of a prefix walk $A_{i}$ not including edge $e$, and a suffix walk $B_{i}$ beginning with edge $e$. We can similarly split $W_{j}$ uniquely
@@
Indeed, for all $l\in[r]$ with $l<i$, the walk $W^{\prime}_{l}=W_{l}$ does not intersect any other walk in $\mathcal{C}$ at an edge, by the definition of $i$. Since the multiset of edges traversed by walks in $\mathcal{C}\setminus\left\{W_{l}\right\}$ and $\mathcal{C}\setminus\left\{W^{\prime}_{l}\right\}$ are the same, this means that $W^{\prime}_{l}$ does not intersect any other walk in $\mathcal{C}^{\prime}$ at an edge either. So $i$ is the smallest index in $[r]$ such that $W^{\prime}_{i}$ intersects some other walk in $\mathcal{C}^{\prime}$ at an edge. Since the prefixes of $W^{\prime}_{i}$ and $W_{i}$ before edge $e$ are the same, we see that $e$ is also the first edge in $W^{\prime}_{i}$ which is contained in another walk of $\mathcal{C}^{\prime}$. Then because $W_{j}^{\prime}$ traverses edge $e$, and $W_{l}^{\prime}=W_{l}$ for all $l\not\in\left\{i,j\right\}$, we get that $j>i$ is the smallest index such that $W_{j}^{\prime}$ overlaps with $W^{\prime}_{i}$ at edge $e$. Then, when we swap the suffixes of $W^{\prime}_{i}$ and $W^{\prime}_{j}$ after the first appearance of $e$ on these walks, we recover $W_{i}$ and $W_{j}$ respectively, and so applying the suffix swapping procedure to $\mathcal{C}^{\prime}$ produces the original collection $\mathcal{C}$ as claimed.
@@
$\xi(\mathcal{C})=\xi(\mathcal{C}^{\prime})$
@@
$\xi(\mathcal{C})+\xi(\mathcal{C}^{\prime})=0.$
@@
$\sum_{\mathcal{C}\in\mathcal{S}}\xi(\mathcal{C})=0.$
@@
In the proof of Lemma 4.7, our argument used the fact that we work over a field of characteristic two. This restriction on the characteristic is only included for the sake of simplicity, and is not necessary for Lemma 4.7 to hold. Indeed, if we instead worked over a field of odd characteristic, then all that changes is that terms in the expansion of the determinant come with a sign, and we can now pair up and cancel terms with opposite signs.
@@
$\det\leavevmode\nobreak\ (I-X)^{-1}[S,T]=\sum_{\ell=1}^{\infty}\left(\sum_{\mathcal{C}\in\mathcal{D}_{\ell}(S,T)}\xi(\mathcal{C})\right).$
@@
Suppose $\mathcal{C}=\langle P_{1},\dots,P_{r}\rangle$ is a collection of $r$ edge-disjoint paths from $S$ to $T$ in $G$. Then the term $\xi(\mathcal{C})$ occurs in the expansion of
@@
given by Corollary 4.10. Moreover, any collection of paths $\mathcal{C}^{\prime}\neq\mathcal{C}$ from $S$ to $T$ has weight $\xi(\mathcal{C}^{\prime})\neq\xi(\mathcal{C})$, because $\mathcal{C}$ consists of edge-disjoint paths (so looking at the variables appearing in $\xi(\mathcal{C})$, we can recover $\mathcal{C}$ uniquely). Hence, no other term from the summation in Corollary 4.10 produces the same monomial $\xi(\mathcal{C})$. So $\xi(\mathcal{C})$ appears in Equation 10 with nonzero coefficient, which implies that the determinant from Equation 10 is a nonzero formal power series.
@@
DAGs are an interesting case for APC and $k$-APC, since we do not know of algorithms solving these problems on DAGs faster than the case of general directed graphs, and the best conditional lower bounds we have for these problems hold in the case of DAGs [AGI+19].
@@
First, any walk in a DAG has length at most $n-1$. Hence over DAGs, $X^{\ell}$ is the all-zeros matrix for $\ell\geq n$, so Proposition 3.3 shows that $\Gamma=(I-X)^{-1}$ is a matrix whose entries are polynomials. Thus, determinants of submatrices of $\Gamma$ are just polynomials, instead of formal power series involving infinite sums. Second, any walk in a DAG is a path. Thus we can skip the step in Lemma 4.12 of going from edge-disjoint walks to edge-disjoint paths.